OpenStudy (codyhelton):

The least common multiple of 20, 24, and 45 is _____. a) 30 b) 180 c) 360 d) 21,600

OpenStudy (mariamikayla):

20|2 10|2 5|5 1 20=2^2 *5 24|2 12|2 6|2 3|3 1 24=2^3 *3 45|3 15|3 5|5 1 45=3^2 * 5 the least common multiple is : 2^3 * 3^2 *5= 8*9*5= 40*9= 360